OpenStudy (anonymous):

3-n/2=3/4 It's a fractional coefficient, and i keep getting -3/2 yet that's not the right answer apparently, not sure if I'm doing it wrong or not.

OpenStudy (amistre64):

(3-n)/2 or 3 - (n/2) ??

OpenStudy (amistre64):

\[\frac{3-n}{2}=\frac{3}{4}\] \[\frac{4(3-n)}{2}=\frac{4(3)}{4}\] \[2(3-n)=3\] \[6-2n=3\] \[-2n=-3~:~n=\frac32\] or ......

OpenStudy (amistre64):

\[3-\frac{n}{2}=\frac{3}{4}\] \[4(3)-\frac{4n}{2}=\frac{4(3)}{4}\] \[12-2n=3\] \[-2n=-9~:~n=\frac92\]

OpenStudy (anonymous):

Thanks

OpenStudy (amistre64):

youre welcome